# HyperSIGMA: Hyperspectral Intelligence Comprehension Foundation Model

### v1

@@ Abstract
Foundation models (FMs) are revolutionizing the analysis and understanding of remote sensing (RS) scenes, including aerial RGB, multispectral, and SAR images. However, hyperspectral images (HSIs), which are rich in spectral information, have not seen much application of FMs, with existing methods often restricted to specific tasks and lacking generality. To fill this gap, we introduce HyperSIGMA, a vision transformer-based foundation model for HSI interpretation, scalable to over a billion parameters. To tackle the spectral and spatial redundancy challenges in HSIs, we introduce a novel sparse sampling attention (SSA) mechanism, which effectively promotes the learning of diverse contextual features and serves as the basic block of HyperSIGMA. HyperSIGMA integrates spatial and spectral features using a specially designed spectral enhancement module. In addition, we construct a large-scale hyperspectral dataset, HyperGlobal-450K, for pre-training, which contains about 450K hyperspectral images, significantly surpassing existing datasets in scale. Extensive experiments on various high-level and low-level HSI tasks demonstrate HyperSIGMA’s versatility and superior representational capability compared to current state-of-the-art methods. Moreover, HyperSIGMA shows significant advantages in scalability, robustness, cross-modal transferring capability, and real-world applicability. The code and models will be released at HyperSIGMA.
@@ 1 Introduction
Figure 1: HyperSIGMA offers a universal solution for HSI processing, demonstrating superior performance across 16 datasets and 7 tasks, including both high-level and low-level hyperspectral tasks, as well as multispectral scenes. It outperforms advanced models like SpectralGPT, even those specifically designed for these tasks.
@@
The challenges of processing HSIs primarily stem from their inherent characteristics: high dimensionality, data redundancy, and spatial variability [11]. High dimensionality often leads to the Hughes phenomenon [12], where machine learning models overfit as the number of channels increases, especially with limited training samples. Data redundancy, both spectral and spatial, results in unnecessary computations due to uninformative bands and pixels, often linked to the sensors’ spectral and spatial resolutions. Spatial variability, influenced by imaging conditions like atmosphere, lighting, and topography, causes mismatches between object categories and spectral curves. To address these challenges, traditional HSI interpretation strategies typically involve two steps: dimensionality reduction [13] and feature extraction [14], with feature extraction being a critical research focus due to its impact on subsequent tasks’ performance. The HSI community continually seeks effective feature extractors, leading to numerous proposed methods [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Among these, deep learning techniques have become predominant for their ability to automatically extract strong feature representations. To tackle spatial variability and enhance contextual feature representation, significant progress has been made with HSI interpretation methods based on convolutional neural networks (CNNs) [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38], recurrent neural networks (RNNs) [39, 40, 41, 42, 43, 44, 45, 46], transformers [47, 48, 49, 50, 51, 52, 53, 54, 55, 56], and state space models [57, 58, 59, 60]. Despite these advancements, most HSI methods remain scene-independent, meaning models are trained and tested on the same specialized scenarios with limited cross-scene knowledge transfer, even for similar tasks. For instance, models for classifying Indian Pines and Pavia University scenes are trained separately, restricting their applicability. Therefore, developing universal HSI processing methods is crucial for advancing the hyperspectral community. Fig. 2 illustrates the differences between traditional and universal HSI processing schemes.
Foundation models (FMs), an emerging technology rich in universal knowledge encoding, offer a promising solution to this issue. Recently, FMs have revolutionized artificial intelligence, attracting widespread attention for their exceptional performance across diverse tasks and their potential to achieve artificial general intelligence [61]. As defined in [62], FMs are built by training on massive datasets, optimizing numerous parameters with extensive computing power, which enables them to generalize well across different scenarios by leveraging the learned universal knowledge. Most FMs utilize the transformer architecture [63], known for its scalability and flexibility, allowing for rapid parameter scaling through the stacking of transformer blocks. FMs first saw success in natural language processing (NLP) with models like BERT [64], XLNet [65], and the GPT series [66, 67, 68, 69], including ChatGPT. This success has inspired the creation of large-scale vision foundation models [70, 71, 72, 73, 74, 75] based on vision transformers (ViT) [76], which have demonstrated impressive performance in various visual tasks.
@@
Remote sensing images, as a distinct type of visual data, have conventionally been interpreted by directly fine-tuning existing models [77, 78], typically pre-trained on natural image datasets such as ImageNet [79]. However, this approach may be suboptimal due to the domain gap between remote sensing (RS) and natural images [80]. Moreover, RS images, abundant in diverse resolutions, temporal variations, wavelengths, and modalities, are widely accessible on a large scale through continuous satellite monitoring. Consequently, recent studies have introduced ViT-based [76] RS foundational models pre-trained on RS data. Existing RS foundation models primarily target high-resolution aerial RGB images [81, 82, 80, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96], multispectral images [97, 3], SAR images [98, 99], or multimodal images [100, 101, 102, 103, 104, 105, 106, 107]. However, there are few FMs specifically tailored for HSI interpretation. This gap may be attributed to the unique challenges of HSIs in data collection, large-scale pre-training, and model structure design.
While existing RS foundation models can be applied to HSI interpretation tasks, they are hindered by domain gaps. Fig. 3 visualizes the image samples, acquisition sensors, and platforms of different data types. RGB images of natural scenes, captured by cameras, consist of red, green, and blue bands. SAR images, acquired by UAVs or satellites with SAR cameras, typically have two or four bands. Multispectral images, obtained by UAVs or satellites with multispectral cameras, often include red, green, blue, and near-infrared bands. Commonly used aerial RGB images, often sourced from online databases like Google Earth, are derived from corresponding channels of multispectral RS images. In contrast, HSIs are captured by UAVs or satellites using hyperspectral cameras and contain hundreds of bands, allowing them to depict terrestrial objects at a finer wavelength scale compared to aerial RGB, SAR, and multispectral images.
@@
In this paper, we present HyperSIGMA (HyperSpectral IntelliGence coMprehension foundAtion model), the first foundation model specifically designed for HSI interpretation. HyperSIGMA integrates spatial and spectral features using a specially designed spectral enhancement module. To tackle the spectral and spatial redundancy challenges in HSIs, we introduce a novel sparse sampling attention (SSA) mechanism, which effectively promotes the learning of diverse contextual features and serves as the foundational block of HyperSIGMA. In addition, we have constructed a large-scale hyperspectral dataset, HyperGlobal-450K (Hyperspectral Global Image dataset), for pre-training. This dataset comprises about 450K hyperspectral images, equivalent to over 20 million trispectral images with non-overlapping channels. Drawing inspiration from the scaling law [108] and the success of existing large-scale RS foundation models [87, 3, 100], we have scaled HyperSIGMA to over 1 billion parameters, supported by HyperGlobal-450K. Unlike existing RS foundation models focusing primarily on high-level tasks such as semantic segmentation and object detection, with minimal exploration of low-level tasks like image denoising and super-resolution, HyperSIGMA offers a unified solution to both high-level and low-level tasks (see Fig. 1). We hope this study provides valuable insights into developing FMs for HSIs and anticipate that HyperSIGMA’s strong representation capability will advance the hyperspectral RS field across diverse applications.
@@
  • Extensive experiments across diverse HSI tasks provide valuable insights into HyperSIGMA’s remarkable versatility and superior representational capability compared to current state-of-the-art methods. Moreover, HyperSIGMA demonstrates significant advantages in scalability, robustness, cross-modal transferring capability, and real-world applicability.
@@ 2 Related Work
2.1 Hyperspectral Image Processing
Traditional HSI dimensionality reduction approaches primarily fall into two categories: band selection and channel compression. Band selection aims to identify the most discriminative channels by employing various criteria [13, 109, 110]. On the other hand, channel compression techniques project valuable information into a lower-dimensional space. Examples of such methods include principal component analysis (PCA) [111], independent component analysis (ICA) [112], and linear discriminant analysis (LDA) [113]. While these methods have been pivotal in traditional HSI processing, which typically processes a single image, their application to datasets like HyperGlobal-450K poses challenges due to the massive number of samples. Utilizing these methods in such contexts not only restricts data diversity but also introduces considerable computational overhead. To address this, we adopt a strategy during pre-training where, before feeding the data into the model, we perform dimensionality reduction by randomly selecting a fixed number of channels for each batch of samples in every iteration. This approach significantly enhances data richness while ensuring training efficiency.
Early HSI feature extraction methods relied on spectral data, either in raw pixel form or after dimensionality reduction, for various machine learning algorithms like k-nearest neighbors [114], multinomial logistic regression [115], random forest [116], and support vector machine [117]. Spatial information was later incorporated through handcrafted descriptors such as morphological profiles [118], superpixels [119], gray-level co-occurrence matrices [120], and filter algorithms [121]. However, these methods relied heavily on expert knowledge and offered limited performance [122]. Nowadays, deep learning-based extractors, particularly CNNs, emerge as the mainstream solution, capturing spectral [26], spatial [123], and spectral-spatial features [29] at various levels of abstraction [124, 125, 126, 127, 128, 129, 130, 131]. CNNs gradually replaced fully connected networks [132, 133] due to their ability to preserve data structure and local perception. Attention mechanisms further enhanced feature extraction by assigning weights to channels [134] or spatial positions [135], improving long-range perception [63, 136, 137]. Recently, transformer-based networks [48, 51] have become popular for their superior context modeling capability, leveraging Multi-Head Self-Attention (MHSA). Nevertheless, existing extractors lacked generalization across scenarios since the models are trained separately on each scene. In this paper, we propose HyperSIGMA, which leverages pre-training on large-scale hyperspectral data for generalized feature extraction.
2.2 Remote Sensing Foundation Model
Existing RS foundation models are either supervised or unsupervised, depending on the pre-training method. [80] pretrains CNNs and ViT models using the Million-AID dataset [138], which contains a million labeled RS scene-level images. However, due to the high cost of annotating RS datasets and the abundance of unlabeled images, unsupervised pre-training via self-supervised learning is more prevalent. Among them, contrastive learning [139, 140, 141] leverages RS-specific characteristics such as seasonal [81], temporal [142], and geographic variations [143]. However, these methods require carefully designed pretext tasks, which can be inefficient to prepare. To address this, many models [83, 84, 97, 88, 85, 87, 91, 94, 95, 3, 102] employ masked image modeling (MIM) [144, 145, 146], which only requires large unlabeled datasets. For instance, SatMAE [97] extends MAE [146] to multispectral and multitemporal data. SpectralGPT [3] partitions multispectral images into 3-D cubes to create visual embeddings for MAE. DOFA [102], inspired by neural plasticity, learns a projection matrix to generate visual embeddings based on input wavelengths. However, these approaches primarily focus on aerial RGB and multispectral images. In contrast, our HyperSIGMA model, with over 1 billion parameters, is specifically designed for HSI interpretation and pre-trained on the large-scale HyperGlobal-450K HSI dataset using MAE [146], making it the world’s first billion-level HSI foundation model.
2.3 Large-Scale Remote Sensing Dataset
Pre-training of RS foundation models typically relies on large-scale datasets. For aerial RGB images, the fMoW-RGB [147] dataset supports the development of numerous classical RS models [82, 97, 88, 95, 94]. Similarly, the MillionAID dataset [138], comparable to ImageNet-1K [79] with one million samples, is widely utilized [80, 84, 87, 89, 90]. Recent advancements in RS multitask pre-training [93] and the RS Vision-Language model GeoChat [148] leverage the SAMRS dataset [92], a large-scale RS dataset with instance and segmentation labels built upon existing RS object detection datasets [149, 150, 151, 152], containing 105,090 images and 1,668,241 instances, far surpassing the capacity of existing high-resolution RS segmentation datasets [153, 154]. For multispectral images, Cong et al. [97] introduced a multispectral version of fMoW-RGB using Sentinel-2 images, called fMoW-Sentinel. Hong et al. [3] further selected 1,067,070 Sentinel-2 images from fMoW-Sentinel and BigEarthNet-S2 [155] to pre-train the SpectralGPT model. Besides, multimodal models [156, 98, 101] that incorporate SAR images are pre-trained on the SSL4EO-S12 dataset [157]. Other datasets like GeoPile [91], Satlas [158], and SeCo-1M [81] also contribute significantly to advancing RS foundation models. Notably, our HyperGlobal-450K stands out as the first large-scale hyperspectral pre-training dataset, containing about 450K hyperspectral images. This volume equates to over 20 million trispectral images with non-overlapping channels, facilitating the development of hyperspectral foundation models.
Figure 4: The distribution of HyperGlobal-450K samples across the globe, comprising 1,701 images (1,486 EO-1 and 215 GF-5B) with hundreds of spectral bands.
2.4 Self-Attention Mechanism
The MHSA module serves as the cornerstone of transformer [63, 76], enabling adaptive global context capture. However, its computational complexity scales quadratically with input token length. Swin Transformer [70] addresses this issue by introducing WMHSA, which confines attention within non-overlapping windows, leading to linear complexity. Nevertheless, fixed-size windows in WMHSA struggle with diverse object sizes and distributions in real scenes. VSA [159] addresses this by introducing adaptive scaling and offsetting of windows, later extended to quadrangles [160], enhancing attention flexibility. RVSA [84] builds upon VSA, introducing learnable rotation for objects with varied orientations, and enhancing context extraction. Besides window-based approaches, deformable mechanisms have gained traction. Deformable convolution [161] adapts convolution kernel sampling positions, followed by attention methods like Deformable Attention [162] and DMHA [163]. In this paper, we propose a novel sparse sampling attention mechanism that enhances flexibility by adaptively capturing contexts within deformable regions, each with a few sampling points to handle redundancy in HSIs.
@@ 3.1 Data Source
[…]
@@ 3.2 Data Acquisition and Processing
We acquired the data by downloading it from official websites. For the EO-1 Hyperion data, we selected images from 2011 to 2017 to ensure full coverage of global locations (Path/Row). We chose data with less than 5% cloud cover and excluded those with unknown cloud cover percentages. For the GF-5B data, we selected 152 images from Nei Mongol and Hebei Province, 24 images from Henan Province, and 39 images from Nei Mongol, Jilin Province, and Liaoning Province in China, as shown in Fig. 4. By combining EO-1 and GF-5B data, the HyperGlobal-450K dataset spans all continents, comprising 1,486 scenes of 2,000$\times$256 EO-1 HSIs and 215 scenes of 2,000$\times$2,000 GF-5B images. After removing bad bands and water vapor absorption bands, the EO-1 HSIs contain 175 bands. For the GF-5B HSIs, considering typical RS land object recognition, we used 150 spectral bands ranging from 0.4 to 1.0 $\mu$m (visible light and near-infrared). All collected HSIs were cut into non-overlapping patches of size 64$\times$64 for model pre-training. This resulted in 447,072 HSI patches, including 247,072 EO-1 patches and 200,000 GF-5 patches, providing rich spatial and spectral characteristics. Fig. 4 shows sample patches of typical landscapes from different regions, including forests, grasslands, barelands, and croplands, which clearly exhibit the characteristics of their respective geographical regions. For instance, patches from tropical rainforests in South America (Region B) and glaciers in Antarctica (Region D) are distinctly green and white, demonstrating the diversity of HyperGlobal-450K. In summary, HyperGlobal-450K, with its global geographic images, provides valuable data for subsequent pre-training.

### v2

@@ Abstract
Accurate hyperspectral image (HSI) interpretation is critical for providing valuable insights into various earth observation-related applications such as urban planning, precision agriculture, and environmental monitoring. However, existing HSI processing methods are predominantly task-specific and scene-dependent, which severely limits their ability to transfer knowledge across tasks and scenes, thereby reducing the practicality in real-world applications. To address these challenges, we present HyperSIGMA, a vision transformer-based foundation model that unifies HSI interpretation across tasks and scenes, scalable to over one billion parameters. To overcome the spectral and spatial redundancy inherent in HSIs, we introduce a novel sparse sampling attention (SSA) mechanism, which effectively promotes the learning of diverse contextual features and serves as the basic block of HyperSIGMA. HyperSIGMA integrates spatial and spectral features using a specially designed spectral enhancement module. In addition, we construct a large-scale hyperspectral dataset, HyperGlobal-450K, for pre-training, which contains about 450K hyperspectral images, significantly surpassing existing datasets in scale. Extensive experiments on various high-level and low-level HSI tasks demonstrate HyperSIGMA’s versatility and superior representational capability compared to current state-of-the-art methods. Moreover, HyperSIGMA shows significant advantages in scalability, robustness, cross-modal transferring capability, real-world applicability, and computational efficiency. The code and models will be released at HyperSIGMA.
@@ 1 Introduction
Figure 1: HyperSIGMA offers a universal solution for HSI processing, demonstrating superior performance across 20 datasets, including both high-level and low-level hyperspectral tasks, as well as multispectral scenes. It outperforms advanced models like SpectralGPT, even those specifically designed for these tasks. HIC: Hyperspectral Image Classification. HTD: Hyperspectral Target Detection. HAD: Hyperspectral Anomaly Detection. HCD: Hyperspectral Change Detection. HIU: Hyperspectral Image Unmixing. HID: Hyperspectral Image Denoising. HSR: Hyperspectral Super-Resolution. MCD: Multispectral Change Detection.
@@
The challenges of processing HSIs primarily stem from their inherent characteristics: high dimensionality, data redundancy, and spatial variability [11]. High dimensionality often leads to the Hughes phenomenon [12], where machine learning models overfit as the number of channels increases, especially with limited training samples. Data redundancy, both spectral and spatial, results in unnecessary computations due to uninformative bands and pixels, often linked to the sensors’ spectral and spatial resolutions. Spatial variability, influenced by imaging conditions like atmosphere, lighting, and topography, causes mismatches between object categories and spectral curves. To address these challenges, traditional HSI interpretation strategies typically involve two steps: dimensionality reduction [13] and feature extraction [14], with feature extraction being a critical research focus due to its impact on subsequent tasks’ performance. The HSI community continually seeks effective feature extractors, leading to numerous proposed methods [15, 16, 17]. Among these, deep learning techniques have become predominant for their ability to automatically extract strong feature representations. To tackle spatial variability and enhance contextual feature representation, significant progress has been made with HSI interpretation methods based on convolutional neural networks (CNNs) [18], recurrent neural networks (RNNs) [19], transformers [20], and state space models [21] [^1] The development history of HSI processing technology can be found in the appendix.. Despite these advancements, most HSI methods remain scene-dependent, meaning models are trained and tested on the same specialized scenarios with limited cross-scene knowledge transfer, even for similar tasks. For instance, models for classifying Indian Pines and Pavia University scenes are trained separately, restricting their applicability. Therefore, developing universal HSI processing methods is crucial for advancing the hyperspectral community. Fig. 2 illustrates the differences between traditional and universal HSI processing schemes.
Foundation models, developed through training on vast datasets and optimized using substantial computational resources, have recently been transforming the field of artificial intelligence [22, 23]. Most foundation models utilize the transformer architecture [24], known for its scalability and flexibility, allowing for rapid parameter scaling through the stacking of transformer blocks. In the field of computer vision (CV), numerous large-scale foundation models [25, 26, 27] based on vision transformers  [28, 29] have demonstrated exceptional performance across various tasks. Similarly, current remote sensing (RS) foundation models [30, 31] have proven capable of effectively handling multiple tasks, including scene classification, semantic segmentation, object detection, and change detection.
@@
Intuitively, developing hyperspectral foundation models may provide a promising solution for unified HSI interpretation across tasks and scenes. However, our literature review reveals a lack of foundation models specifically designed for HSI interpretation, despite their ability to capture terrestrial objects at a finer wavelength scale than aerial RGB, SAR, and multispectral images (see Fig. 3). The challenges to developing large-scale foundation models for HSI processing primarily lie in the unique characteristics of hyperspectral data, which complicate data collection, pre-training, and model design. Specifically, large-scale pre-training requires significant computational resources, while acquiring and processing hyperspectral data is labor- and time-intensive. Additionally, effective pre-training strategies and model architectures are expected to be tailored to the distinct characteristics of HSIs. These combined obstacles hinder the progress of large-scale hyperspectral foundation models in the RS community.
@@
To address these challenges, we introduce HyperSIGMA (HyperSpectral IntelliGence coMprehension foundAtion model), the first step towards hyperspectral foundation models tailored for HSI interpretation. HyperSIGMA integrates spatial and spectral features using a specially designed spectral enhancement module. To tackle the issues of spectral and spatial redundancy in HSIs, we introduce a novel sparse sampling attention (SSA) mechanism, which effectively promotes the learning of diverse contextual features and serves as the foundational block of HyperSIGMA. In addition, we have constructed a large-scale hyperspectral dataset, HyperGlobal-450K (Hyperspectral Global Image dataset), for pre-training. This dataset comprises about 450K hyperspectral images, equivalent to over 20 million trispectral images with non-overlapping channels. The appendix presents detailed comparisons between HyperGlobal-450K and existing large-scale RS datasets. Drawing inspiration from the scaling law [32] and the success of existing large-scale RS foundation models [33, 3, 34], we have scaled HyperSIGMA to over 1 billion parameters, supported by HyperGlobal-450K. Unlike existing RS foundation models focusing primarily on high-level tasks such as semantic segmentation and object detection, with minimal exploration of low-level tasks like image denoising and super-resolution, HyperSIGMA offers a unified solution to both high-level and low-level tasks (see Fig. 1). We hope this study provides valuable insights into developing foundation models for HSIs and anticipate that HyperSIGMA’s strong representation capability will advance the hyperspectral RS field across diverse applications.
@@
  • Extensive experiments across diverse HSI tasks provide valuable insights into HyperSIGMA’s remarkable versatility and superior representational capability compared to current state-of-the-art methods. Moreover, HyperSIGMA demonstrates significant advantages in scalability, robustness, cross-modal transferring capability, real-world applicability and computational efficiency.
Figure 4: The distribution of HyperGlobal-450K samples across the globe. The sampled patches of typical landscapes from different regions, including forests, grasslands, barelands, and croplands, clearly exhibit the characteristics of their respective geographical regions.
@@ 2 Related Work
Numerous foundation models have been proposed for natural images, with the vision transformer adopted as the mainstream network structure. Among them, scaling is an effective strategy. Typically, ViT [29] and Swin Transformer [28] have been scaled up to 1 or 10 billion parameters [27, 35] to explore the potential of large-scale vision models. Another available approach is combining the advantages of both convolution and attention. For this purpose, ViTAE [25] introduces the inductive bias of CNN into vision transformers. In addition to vision transformers, CNN-based large-scale foundation models have also been explored [36].
Compared to the large model size in the CV field, the foundation models in the RS field are usually smaller, focusing more on the pre-training strategy on RS images [37]. Due to the high cost of annotating RS datasets and the abundance of unlabeled images, pre-training via self-supervised learning is more prevalent. Among them, contrastive learning [38] leverages RS-specific characteristics such as seasonal [39], temporal [40], and geographic variations [41]. Besides, many models [30, 31, 42, 43, 33, 44, 3] employ masked image modeling (MIM) [45, 46]. For instance, SatMAE [42] extends MAE [46] to multispectral and multitemporal data. SpectralGPT [3] partitions multispectral images into 3-D cubes to create visual embeddings for MAE. While SMLFR [43] develops pioneering generative CNN foundation models for RS scenes by pre-training on approximately 9 million images, achieving superior performance on multiple RS downstream tasks. In addition to pre-training from scratch, some RS foundation models are constructed based on existing vision models. GFM [44] utilizes a powerful universal foundation model to guide the learning of RS features, while MTP [47] performs continuous multi-task supervised pre-training on existing RS or CV foundation models. In contrast, UPetu [48] introduces the first unified RS parameter-efficient fine-tuning framework, achieving performance comparable to full fine-tuning while optimizing just 0.73% of the parameters.
Compared to existing approaches that primarily focus on RGB and multispectral images, HyperSIGMA is the first billion-level foundation model specifically for HSI interpretation. Leveraging pre-training on the large-scale HyperGlobal-450K dataset, HyperSIGMA employs a sparse sampling mechanism to tackle spectral and spatial redundancy in HSIs, offering a unified approach to both high-level and low-level hyperspectral vision tasks. Further detailed comparison between HyperSIGMA and existing representative CV and RS foundation models can be found in the appendix, where more discussions about the related works of HSI processing techniques, MHSA-based vision transformer networks, and large-scale RS datasets are included.
@@ 3.1 Data Source
In view of a series of significant criteria such as global coverage and free access, we select both the Earth Observing One (EO-1) and Gaofen5 (GF-5) satellites as the data source of HyperGlobal-450K. The details of the motivation and workflow in selecting sensors and images, and the basic information of these satellites are presented in the appendix.
@@ 3.2 Data Acquisition and Processing
We collect all EO-1 images acquired during 2011-2017 and additionally compensate the GF-5 images from China. Then, we design a set of standards to achieve image selection from cloud contents, locations, and bands. After clipping, we obtain a large-scale hyperspectral dataset with global coverage (see Fig. 4), containing 447,072 HSI patches in size of 64$\times$64, i.e., HyperGlobal-450K, including 247,072 EO-1 patches and 200,000 GF-5 patches. Detailed data acquisition and processing pipelines can be found in the appendix.